OpenStudy (anonymous):

please someone im taking a test Which of the following is the simplified version of log 3 x + 2 log 3 y + 4 log 3 z ? please someone im taking a test Which of the following is the simplified version of log 3 x + 2 log 3 y + 4 log 3 z ? @Mathematics

OpenStudy (anonymous):

Should you not list the 'following' options?

OpenStudy (anonymous):

\[\log_{3} \sqrt{xy^2/z^4}\]

OpenStudy (anonymous):

or

OpenStudy (anonymous):

\[\log_{3} \sqrt{x}/y^2z^4\]

OpenStudy (anonymous):

or

OpenStudy (anonymous):

The numbers you list after log is a numerator or base? ie is it all log to the base 3?

OpenStudy (anonymous):

numerator and then where u see / starts the denomintor

OpenStudy (anonymous):

\[\log_{3} \sqrt{x}y^2z^4\]

OpenStudy (anonymous):

or

OpenStudy (anonymous):

\[\log_{3} \sqrt{xz^4/y^2}\]

OpenStudy (anonymous):

I got \[\log_{3} xy^2z^4\] But I'm not too sure of your notation, every answer seems to have a square root in it but I can't see why.

OpenStudy (anonymous):

thank u it was right!

OpenStudy (anonymous):

How was it right? It wasn't even one of the options?

OpenStudy (anonymous):

it was lol i just couldnt get it wrote down properly!

OpenStudy (anonymous):

can u help me with this one? What is the solution to the equation log 3 x – log 3 2 = 2 ?

OpenStudy (anonymous):

You understand how I got it or want me to work through it before your test?

OpenStudy (anonymous):

im in the test:(

OpenStudy (anonymous):

\[\log_{3} x-\log_{3} 2=2\]

OpenStudy (anonymous):

\[\log_{3} (x/2)=2 <=> x/2= 3^2, x/2 = 9, x=9/2\]

OpenStudy (anonymous):

But I'd double check that

OpenStudy (anonymous):

thanks i wish i can give another medal!lol

OpenStudy (anonymous):

If you can for some reason use outside help in the test why not use wolfram?

OpenStudy (anonymous):

not alway accurate:( i just posted another question wana help me answer?

OpenStudy (anonymous):

Ah I messed that one up x/2=9, x=2(9) x=18

OpenStudy (anonymous):

If you can change your last answer I would, it is 18 not 9/2 made an error